### v1

# Two-grid convergence theory for symmetric positive semidefinite linear systems

Abstract.
This paper is devoted to the convergence theory of two-grid methods for symmetric positive semidefinite linear systems, with particular focus on the singular case. In the case where the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver, we derive a succinct identity for characterizing the convergence factor of two-grid methods. More generally, we present some convergence estimates for two-grid methods with approximate coarse solvers, including both linear and general cases. A key feature of our analysis is that it does not require any additional assumptions on the system matrix, especially on its null space.
Key words and phrases:
two-grid methods, symmetric positive semidefinite matrices, singular systems, convergence analysis
2020 Mathematics Subject Classification:
Primary 65F08, 65F10, 65N55; Secondary 15A10, 15A18
1. Introduction
As a powerful solver with linear or near-linear computational complexity, multigrid has been widely used for the numerical solution of partial differential equations; see, e.g., [11, 4, 26, 27, 31]. The foundation of multigrid is a two-grid scheme, which consists of two complementary error-reduction processes: smoothing (or relaxation) and coarse-grid correction. In the correction step, a coarse-grid system with fewer unknowns needs to be solved. In practice, it is often too costly to solve the coarse-grid system exactly. A typical strategy is to apply the two-grid scheme recursively in the correction steps, which will yield an algorithm with multilevel structure (i.e., multigrid method). For symmetric positive definite (SPD) linear systems, the convergence properties of two-grid and multigrid methods have been well studied; see, e.g., [29, 35, 30, 7, 8, 17, 31, 32, 33, 34].
Many systems arising in mathematical models may only be semidefinite, such as discretized partial differential equations with Neumann boundary conditions (see, e.g., [1]) and Laplacian matrices of graphs (see, e.g., [24, 16]). In such cases, the system matrix, denoted by $A$, is singular and symmetric positive semidefinite (SPSD). Some convergence properties of classical iterative methods for SPSD systems can be found, e.g., in [13, 5, 14, 6, 9, 28, 10]. The convergence of Schwarz methods for SPSD systems was studied in [21]. Our focus here is on the convergence of two-grid methods for SPSD systems. In practice, two-grid convergence theory can be used to guide or motivate the design of multigrid methods; see, e.g., [3, 19, 18]. Moreover, two-grid convergence estimates can be applied to the analysis of multigrid methods; see, e.g., [33, 34].
In the SPD case, an identity for the convergence factor of two-grid methods was established in [8]; see [30, 36] for an abstract version. Ideally, one can derive a two-grid convergence identity for the semidefinite case without imposing any restrictions on the null space of $A$, which is denoted by $\mathcal{N}(A)$. Under several essential assumptions on local null spaces, a sharp convergence estimate for the successive subspace correction methods [29, 30] applied to SPSD systems was presented in [15]; see [28] for a simplified analysis. More specifically, [15] offers an identity for the seminorm of error propagation operator, despite it involves a ‘sup-inf-inf’ expression relating to the global null space and its orthogonal complement. In [2], a classical two-grid convergence estimate for SPD systems [20, 23, 25] was extended to the semidefinite case. A two-grid convergence theory for singular systems was developed in [22], in which a restrictive condition on the coarse space and $\mathcal{N}(A)$ is required. Even in the special SPSD case, the two-grid estimates in [22] involve an extended prolongation operator, whose construction is based on $\mathcal{N}(A)$. A noteworthy contribution is the two-grid convergence identity appearing in [31, Theorem 5.3], which was established under the condition that $\mathcal{N}(A)$ is contained in the coarse space.
The aforementioned two-grid convergence theory and the associated assumptions rely largely on $\mathcal{N}(A)$, while the properties of $\mathcal{N}(A)$ may not be easily available. Consequently, it might be tough to apply those results or check the required conditions. Motivated by this observation, we establish a new convergence theory for two-grid methods applied to SPSD systems, which lifts the restrictions on $\mathcal{N}(A)$. Our main contribution is twofold.
  • Firstly, we derive a succinct identity for the convergence factor of two-grid methods, provided that the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver (it is called the exact coarse solver).
  • Secondly, we present some convergence estimates for two-grid methods with inexact coarse solvers, including both linear and general cases.
It is worth mentioning that our theory has several distinctive features in comparison with the existing ones, as described below.
  • It does not require any additional assumptions on $A$, especially on $\mathcal{N}(A)$.
  • It yields an identity for the convergence factor of exact two-grid methods instead of an upper bound.
  • It offers some convergence bounds for two-grid methods with inexact coarse solvers.
The rest of this paper is organized as follows. In Section 2, we give a sufficient and necessary condition for two-grid convergence factor to be less than $1$. In Section 3, we present a concise identity for the convergence factor of exact two-grid methods. In Section 4, we study the convergence of two-grid methods with approximate coarse solvers. In Section 5, we make some concluding remarks.
2. Preliminaries
We start with some notation that will be used in the subsequent discussions.
  • For two symmetric matrices $S_{1}$ and $S_{2}$, the symbol $S_{1}\succ S_{2}$ (resp., $S_{1}\succeq S_{2}$) means that $S_{1}-S_{2}$ is positive definite (resp., positive semidefinite).
  • $I_{n}$ denotes the $n\times n$ identity matrix (or $I$ when its size is clear from context).
  • $\mathcal{R}(\cdot)$ denotes the range (or column space) of a matrix.
  • $\mathcal{N}(\cdot)$ denotes the null space of a matrix.
  • $\lambda_{\min}(\cdot)$ and $\lambda_{\max}(\cdot)$ denote the smallest and largest eigenvalues of a matrix, respectively.
  • $\lambda_{i}(\cdot)$ denotes the $i$th smallest eigenvalue of a matrix.
  • $\lambda(\cdot)$ denotes the spectrum of a matrix.
  • $\|\cdot\|_{2}$ denotes the Euclidean norm of a vector.
  • $(\cdot)^{\dagger}$ denotes the Moore–Penrose inverse of a matrix.
For any $A\in\mathbb{R}^{n\times n}\backslash\{0\}$ and $A\succeq 0$, we define
$\|\mathbf{v}\|_{A}:=\sqrt{\mathbf{v}^{T}A\mathbf{v}}\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
and
(2.1) $\|X\|_{A}:=\sup_{\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)}\frac{\|X\mathbf{v}\|_{A}}{\|\mathbf{v}\|_{A}}\quad\forall\,X\in\mathbb{R}^{n\times n}.$
It is easy to check that the function $\|\cdot\|_{A}:\mathbb{R}^{n\times n}\rightarrow[0,+\infty]$ satisfies the following properties:
  • (a)
    $\|X\|_{A}\geq 0$ for all $X\in\mathbb{R}^{n\times n}$;
  • (b)
    $\|\alpha X\|_{A}=|\alpha|\|X\|_{A}$ for all $\alpha\in\mathbb{R}$ and $X\in\mathbb{R}^{n\times n}$;
  • (c)
    $\|X_{1}+X_{2}\|_{A}\leq\|X_{1}\|_{A}+\|X_{2}\|_{A}$ for all $X_{1},X_{2}\in\mathbb{R}^{n\times n}$.
That is, (2.1) defines a seminorm on $\mathbb{R}^{n\times n}$.
The following lemma gives two simple but useful properties of the $A$-seminorm.
Lemma 2.1.
If $X\in\mathbb{R}^{n\times n}$ and $A^{\frac{1}{2}}X=YA^{\frac{1}{2}}$ for some $Y\in\mathbb{R}^{n\times n}$, then
(2.2) $\|X\mathbf{v}\|_{A}\leq\|X\|_{A}\|\mathbf{v}\|_{A}\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
and
(2.3) $\|X\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}.$
Proof.
Note that
$\mathcal{N}(A)=\mathcal{N}\big{(}A^{\frac{1}{2}}\big{)}\quad\text{and}\quad\mathcal{R}(A)=\mathcal{R}\big{(}A^{\frac{1}{2}}\big{)}.$
It can be readily seen from (2.1) that (2.2) holds for $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)$. If $\mathbf{v}\in\mathcal{N}(A)$, then
$\|X\mathbf{v}\|_{A}=\big{\|}A^{\frac{1}{2}}X\mathbf{v}\big{\|}_{2}=\big{\|}YA^{\frac{1}{2}}\mathbf{v}\big{\|}_{2}=0,$
where we have used the condition $A^{\frac{1}{2}}X=YA^{\frac{1}{2}}$. Hence, the inequality (2.2) holds for all $\mathbf{v}\in\mathbb{R}^{n}$.
From (2.1), we have
$\|X\|_{A}$ $=\sup_{\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)}\frac{\big{\|}A^{\frac{1}{2}}X\mathbf{v}\big{\|}_{2}}{\big{\|}A^{\frac{1}{2}}\mathbf{v}\big{\|}_{2}}=\sup_{\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(A)}\frac{\big{\|}YA^{\frac{1}{2}}\mathbf{v}\big{\|}_{2}}{\big{\|}A^{\frac{1}{2}}\mathbf{v}\big{\|}_{2}}$
$=\sup_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|Y\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}},$
which yields the expression (2.3). ∎
Consider the solution of linear systems of the form
(2.4) $A\mathbf{u}=\mathbf{f},$
where $A\in\mathbb{R}^{n\times n}\backslash\{0\}$, $A\succeq 0$, $\mathbf{u}\in\mathbb{R}^{n}$, and $\mathbf{f}\in\mathcal{R}(A)$. When $A$ is singular (namely, $\mathcal{N}(A)\neq\{0\}$), the system (2.4) has infinitely many solutions for any $\mathbf{f}\in\mathcal{R}(A)$. We aim to solve (2.4) using a two-grid method, the output of which is an approximation to a solution to (2.4). Two very basic assumptions involved in the two-grid analysis are made below.
  • Let $M$ be an $n\times n$ matrix such that $\|I-MA\|_{A}\leq 1$.
  • Let $P$ be an $n\times n_{\rm c}$ prolongation (or interpolation) matrix such that the Galerkin coarse-grid matrix $A_{\rm c}:=P^{T}AP$ is nonzero, where $n_{\rm c}\,(<n)$ is the number of coarse variables.
Remark 2.2.
We remark that $M$ is not necessarily nonsingular. Moreover, measured in $\|\cdot\|_{A}$, the smoothing iteration matrix in Algorithm 1 below (i.e., $I-MA$) is not required to be a strict contraction. Due to
$A^{\frac{1}{2}}(I-MA)=\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}},$
it follows from (2.3) that
$\|I-MA\|_{A}^{2}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}^{2}}{\|\mathbf{v}\|_{2}^{2}}=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
where
(2.5) $\overline{M}:=M+M^{T}-M^{T}AM.$
Thus, $\|I-MA\|_{A}\leq 1$ if and only if $\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}\geq 0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$, or, equivalently, $\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}\geq 0$ for all $\mathbf{v}\in\mathbb{R}^{n}\backslash\{0\}$ (i.e., $A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\succeq 0$), because $\mathbb{R}^{n}$ is the direct sum of $\mathcal{R}(A)$ and $\mathcal{N}(A)$, denoted by $\mathbb{R}^{n}=\mathcal{R}(A)\oplus\mathcal{N}(A)$.
With the above assumptions, a two-grid method for solving (2.4) is described by Algorithm 1, in which the Moore–Penrose inverse $A_{\rm c}^{\dagger}$ is used as a coarse solver. For convenience, such an algorithm is called an exact two-grid method.
Algorithm 1  Exact two-grid method.
1:Smoothing: $\mathbf{u}^{(1)}\leftarrow\mathbf{u}^{(0)}+M\big{(}\mathbf{f}-A\mathbf{u}^{(0)}\big{)}$triangleright $\mathbf{u}^{(0)}\in\mathbb{R}^{n}$ is an initial guess
2:Restriction: $\mathbf{r}_{\rm c}\leftarrow P^{T}\big{(}\mathbf{f}-A\mathbf{u}^{(1)}\big{)}$
3:Coarse-grid correction: $\mathbf{e}_{\rm c}\leftarrow A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$
4:Prolongation: $\mathbf{u}_{\rm TG}\leftarrow\mathbf{u}^{(1)}+P\mathbf{e}_{\rm c}$
Remark 2.3.
The output of Algorithm 1 converges to a particular solution to (2.4) as $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}$ tends to zero. Indeed, $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}=0$ implies that $\mathbf{u}-\mathbf{u}_{\rm TG}\in\mathcal{N}(A)$ and hence $A\mathbf{u}_{\rm TG}=A\mathbf{u}=\mathbf{f}$.
From Algorithm 1, we have
(2.6) $\mathbf{u}-\mathbf{u}_{\rm TG}=E_{\rm TG}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)},$
where
$E_{\rm TG}=\big{(}I-PA_{\rm c}^{\dagger}P^{T}A\big{)}(I-MA).$
Define
(2.7) $\Pi:=A^{\frac{1}{2}}PA_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}.$
It can be easily verified that
$\Pi^{T}=\Pi=\Pi^{2},\quad\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)},\quad\text{and}\quad\mathcal{R}(\Pi)=\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}.$
Thus, $\Pi$ is an $L^{2}$-orthogonal projector along (or parallel to) $\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}$ onto $\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}$. Since
$A^{\frac{1}{2}}E_{\rm TG}=(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}},$
it follows from Lemma 2.1 and (2.6) that
$\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\|E_{\rm TG}\|_{A}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A},$
where $\|E_{\rm TG}\|_{A}$, called the convergence factor of Algorithm 1, is given by
(2.8) $\|E_{\rm TG}\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}}{\|\mathbf{v}\|_{2}}.$
Then
$\|E_{\rm TG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
$=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm TG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
where
(2.9) $F_{\rm TG}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
Hence, it holds that $\|E_{\rm TG}\|_{A}<1$ if and only if $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$.
The following lemma provides a sufficient and necessary condition (namely, (2.10)) for $\|E_{\rm TG}\|_{A}<1$. It is worth pointing out that (2.10) is not really a restriction on $\mathcal{N}(A)$. In view of (2.10), a sufficient condition for $\|E_{\rm TG}\|_{A}<1$ is $\overline{M}\succ 0$, which in fact can be satisfied by the most commonly used weighted Jacobi and Gauss–Seidel smoothers; see Remark 2.6. Throughout this paper, we always assume that (2.10) holds.
Lemma 2.4.
Under the assumptions of Algorithm 1, $\|E_{\rm TG}\|_{A}<1$ if and only if
(2.10) $\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)}=\mathcal{N}(A).$
Proof.
Observe first that
$\mathbf{v}^{T}F_{\rm TG}\mathbf{v}=\underbrace{\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}}_{\geq 0}+\underbrace{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}_{\geq 0}\geq 0\quad\forall\,\mathbf{v}\in\mathbb{R}^{n}$
and
$\mathcal{N}(A)\subseteq\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)}.$
[…]
$0=P^{T}(I-AM)A^{\frac{1}{2}}\mathbf{x}=P^{T}(I-AM)A^{\frac{1}{2}}\mathbf{x}_{1}=P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{x}_{1},$
it follows that
$\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{x}_{1}=0.$
In addition,
$A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{x}_{1}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}(\mathbf{x}_{1}+\mathbf{x}_{2})=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{x}=0.$
Hence,
$\mathbf{x}_{1}^{T}F_{\rm TG}\mathbf{x}_{1}=\mathbf{x}_{1}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{x}_{1}+\mathbf{x}_{1}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{x}_{1}=0,$
which contradicts with the condition $\|E_{\rm TG}\|_{A}<1$, or, equivalently, $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$ for all $\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$.
(2.10)$\Rightarrow\|E_{\rm TG}\|_{A}<1$”: Conversely, if (2.10) holds, then $\|E_{\rm TG}\|_{A}<1$. Otherwise, there must exist a vector $\mathbf{y}\in\mathcal{R}(A)\backslash\{0\}$ such that $\mathbf{y}^{T}F_{\rm TG}\mathbf{y}=0$. Then
$\mathbf{y}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{y}=0\quad\text{and}\quad\mathbf{y}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{y}=0,$
which lead to
$\mathbf{y}\in\mathcal{N}\big{(}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
and
$\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{y}\in\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
Thus,
$\mathbf{y}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)},$
which contradicts with (2.10), because $\mathbf{y}\notin\mathcal{N}(A)$. This completes the proof. ∎
Corollary 2.5.
Under the assumptions of Algorithm 1, if $\overline{M}\succeq 0$ and
(2.11) $\mathcal{N}(\overline{M})\cap\mathcal{R}(A)=\{0\},$
then $\|E_{\rm TG}\|_{A}<1$.
Proof.
For any $\mathbf{v}\in\mathcal{N}\big{(}\overline{M}A^{\frac{1}{2}}\big{)}$, we have
$A^{\frac{1}{2}}\mathbf{v}\in\mathcal{N}(\overline{M})\cap\mathcal{R}(A).$
It then follows from (2.11) that $\mathbf{v}\in\mathcal{N}(A)$. Since $\mathbf{v}$ is arbitrary and
$\mathcal{N}(A)\subseteq\mathcal{N}\big{(}\overline{M}A^{\frac{1}{2}}\big{)},$
we obtain
$\mathcal{N}\big{(}\overline{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A),$
which, together with the positive semidefiniteness of $\overline{M}$, yields
$\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\overline{M}^{\frac{1}{2}}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\overline{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A).$
The desired result then follows from Lemma 2.4. ∎
Remark 2.6.
From Corollary 2.5, we deduce that $\|E_{\rm TG}\|_{A}<1$ if $\overline{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, provided that the diagonal entries of $A$ are all nonzero. Let
$A=D+L+L^{T},$
where $D$ and $L$ denote the diagonal and strictly lower triangular parts of $A$, respectively. Without loss of generality, we assume that $D\succ 0$. If some diagonal entry of $D$ is zero, the corresponding row and column in $A$ are both zero, in which case one may solve a reduced version of (2.4) obtained by deleting the zero row and column. If $M=\omega D^{-1}$ with $0<\omega<\frac{2}{\lambda_{\max}(D^{-1}A)}$, then
$\overline{M}=\omega^{2}D^{-1}\big{(}2\omega^{-1}D-A\big{)}D^{-1}\succ 0.$
Another typical choice is $M=(D+L)^{-1}$, in which case one has
$\overline{M}=(D+L)^{-T}D(D+L)^{-1}\succ 0.$
3. Convergence analysis of Algorithm 1
In this section, we present a succinct identity for characterizing the convergence factor $\|E_{\rm TG}\|_{A}$.
Define
$r:=\operatorname*{rank}(A)\quad\text{and}\quad s:=\operatorname*{rank}(A_{\rm c}).$
Since $A_{\rm c}=P^{T}AP$, it follows that $s\leq r$. In the extreme case $s=r$, we have
$\operatorname*{rank}(\Pi)=\operatorname*{rank}\big{(}A^{\frac{1}{2}}P\big{)}=s=r,$
which, combined with the fact $\mathcal{R}(\Pi)\subseteq\mathcal{R}(A)$, leads to
$\mathcal{R}(\Pi)=\mathcal{R}(A).$
Then
$(I-\Pi)\mathbf{v}=0\quad\forall\,\mathbf{v}\in\mathcal{R}(A),$
which, together with (2.8), yields $\|E_{\rm TG}\|_{A}=0$. Such an extreme case seldom occurs in practice. In what follows, we only consider the nontrivial case $s<r$.
We first prove a technical lemma, which gives a reformulation of the convergence factor $\|E_{\rm TG}\|_{A}$.
Lemma 3.1.
Under the assumptions of Algorithm 1 and the condition (2.10), one has
(3.1) $\|E_{\rm TG}\|_{A}=\sqrt{1-\lambda_{n-r+1}(F_{\rm TG})},$
where $F_{\rm TG}$ is given by (2.9).
Proof.
It can be easily seen from (2.7) and (2.9) that
(3.2) $\mathcal{N}(A)\subseteq\mathcal{N}(F_{\rm TG}).$
For any $\mathbf{v}\in\mathcal{N}(F_{\rm TG})$, we have
$0=\mathbf{v}^{T}F_{\rm TG}\mathbf{v}=\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}+\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v},$
from which we deduce that
$\mathbf{v}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\quad\text{and}\quad\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\in\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
Then
$\mathbf{v}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{N}\big{(}P^{T}(I-AM)A^{\frac{1}{2}}\big{)}=\mathcal{N}(A),$
where we have used the condition (2.10). The arbitrariness of $\mathbf{v}$ implies that
(3.3) $\mathcal{N}(F_{\rm TG})\subseteq\mathcal{N}(A).$
Combining (3.2) and (3.3), we obtain
(3.4) $\mathcal{N}(F_{\rm TG})=\mathcal{N}(A)$
and hence
$\mathcal{R}(F_{\rm TG})=\mathcal{R}(A).$
Since $F_{\rm TG}\succeq 0$ and $\operatorname*{rank}(F_{\rm TG})=r$, there exists an orthogonal matrix $Q_{1}\in\mathbb{R}^{n\times n}$ such that
$Q_{1}^{T}F_{\rm TG}Q_{1}=\begin{pmatrix}\Lambda_{r}&0\\ 0&0\end{pmatrix},$
where $\Lambda_{r}\in\mathbb{R}^{r\times r}$ is diagonal and $\Lambda_{r}\succ 0$. We then have
$\mathcal{R}(A)=\mathcal{R}(F_{\rm TG})=\mathcal{R}\bigg{(}Q_{1}\begin{pmatrix}I_{r}\\ 0\end{pmatrix}\bigg{)}.$
Thus,
$\|E_{\rm TG}\|_{A}^{2}$ $=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm TG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
[…]
$=1-\min_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}\Lambda_{r}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\mathbf{v}_{r}}$
$=1-\lambda_{\min}(\Lambda_{r})$
$=1-\lambda_{n-r+1}(F_{\rm TG}),$
which yields the expression (3.1). ∎
Remark 3.2.
A symmetric two-grid scheme can be obtained by adding the postsmoothing step
$\mathbf{u}_{\rm STG}\leftarrow\mathbf{u}_{\rm TG}+M^{T}(\mathbf{f}-A\mathbf{u}_{\rm TG})$
after the prolongation step of Algorithm 1. The error propagation of the symmetric two-grid scheme can be expressed as
$\mathbf{u}-\mathbf{u}_{\rm STG}=E_{\rm STG}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
with
$E_{\rm STG}=(I-M^{T}A)\big{(}I-PA_{\rm c}^{\dagger}P^{T}A\big{)}(I-MA).$
Since
$A^{\frac{1}{2}}E_{\rm STG}=\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}}=(I-F_{\rm TG})A^{\frac{1}{2}},$
we obtain
$\|E_{\rm STG}\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\|(I-F_{\rm TG})\mathbf{v}\|_{2}}{\|\mathbf{v}\|_{2}}.$
Then
$\|E_{\rm STG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}(I-F_{\rm TG})^{2}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
[…]
$=\max_{\mathbf{v}_{r}\in\mathbb{R}^{r}\backslash\{0\}}\frac{\mathbf{v}_{r}^{T}(I_{r}-\Lambda_{r})^{2}\mathbf{v}_{r}}{\mathbf{v}_{r}^{T}\mathbf{v}_{r}}.$
Due to
$Q_{1}\begin{pmatrix}I_{r}-\Lambda_{r}&0\\ 0&I_{n-r}\end{pmatrix}Q_{1}^{T}=I-F_{\rm TG}\succeq 0\quad\text{and}\quad\Lambda_{r}\succ 0,$
it follows that $\lambda(\Lambda_{r})\subset(0,1]$. Hence,
$\|E_{\rm STG}\|_{A}^{2}=\lambda_{\max}\big{(}(I_{r}-\Lambda_{r})^{2}\big{)}=\big{(}1-\lambda_{\min}(\Lambda_{r})\big{)}^{2}=\|E_{\rm TG}\|_{A}^{4},$
which yields
$\|E_{\rm STG}\|_{A}=\|E_{\rm TG}\|_{A}^{2}.$
This shows that there is no essential difference between the convergence analysis of Algorithm 1 and that of its symmetrized version.
Based on Lemma 3.1, we can derive the following convergence identity, which is the main result of this section.
Theorem 3.3.
Define
(3.5) $\widetilde{M}:=M+M^{T}-MAM^{T}$
and
(3.6) $\Pi_{A}:=PA_{\rm c}^{\dagger}P^{T}A.$
Under the assumptions of Lemma 3.1, it holds that
(3.7) $\|E_{\rm TG}\|_{A}=\sqrt{1-\sigma_{\rm TG}},$
where
(3.8) $\sigma_{\rm TG}=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}.$
Proof.
Due to
$F_{\rm TG}=I-\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)},$
it follows that
$\lambda(F_{\rm TG})$ $=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{)}$
$=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}(I-\Pi)\big{)}$
$=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi\big{)}.$
According to the proof of Lemma 3.1, we deduce that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi$ has $r$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r$.
Since $\Pi$ is an $L^{2}$-orthogonal projector and $\operatorname*{rank}(\Pi)=s$, there exists an orthogonal matrix $Q_{2}\in\mathbb{R}^{n\times n}$ such that
(3.9) $Q_{2}^{T}\Pi Q_{2}=\begin{pmatrix}I_{s}&0\\ 0&0\end{pmatrix}.$
Let
(3.10) $Q_{2}^{T}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}Q_{2}=\begin{pmatrix}X_{1}&X_{2}\\ X_{2}^{T}&X_{3}\end{pmatrix},$
where $X_{1}\in\mathbb{R}^{s\times s}$, $X_{2}\in\mathbb{R}^{s\times(n-s)}$, and $X_{3}\in\mathbb{R}^{(n-s)\times(n-s)}$. Then
$A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi=Q_{2}\begin{pmatrix}I_{s}&X_{2}\\ 0&X_{3}\end{pmatrix}Q_{2}^{T},$
from which we deduce that $X_{3}$ has $r-s$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r$.
Direct computation yields
$\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}=I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}=Q_{2}\begin{pmatrix}I_{s}-X_{1}&-X_{2}\\ -X_{2}^{T}&I_{n-s}-X_{3}\end{pmatrix}Q_{2}^{T}.$
The positive semidefiniteness of $I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}$ implies that $\lambda_{\max}(X_{3})\leq 1$. Thus,
$\lambda_{n-r+1}(F_{\rm TG})=\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)+\Pi\big{)}=\lambda_{n-r+1}(X_{3}).$
Since
$(I-\Pi)A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)=Q_{2}\begin{pmatrix}0&0\\ 0&X_{3}\end{pmatrix}Q_{2}^{T},$
we obtain
$\lambda_{n-r+1}(X_{3})$ $=\lambda_{n-r+s+1}\big{(}(I-\Pi)A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)\big{)}$
$=\lambda_{n-r+s+1}\big{(}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)A^{\frac{1}{2}}\big{)}$
$=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}.$
The identity (3.7) then follows by using (3.1). ∎
The identity (3.7) provides a direct way to analyze the convergence properties of Algorithm 1. A two-sided estimate for $\|E_{\rm TG}\|_{A}$ can be obtained by applying (3.7) and the following eigenvalue inequality; see, e.g., [12, Corollary 4.3.37].
Lemma 3.4.
Let $H\in\mathbb{C}^{n\times n}$ be Hermitian, and let $\{\mathbf{q}_{k}\}_{k=1}^{m}\subset\mathbb{C}^{n}\,(1\leq m\leq n)$ be a set of orthonormal vectors. Then, for any $i=1,\ldots,m$, it holds that
$\lambda_{i}(H)\leq\lambda_{i}(\check{H})\leq\lambda_{i+n-m}(H),$
where $\check{H}=\big{(}\mathbf{q}_{i}^{\ast}H\mathbf{q}_{j}\big{)}\in\mathbb{C}^{m\times m}$ and $(\cdot)^{\ast}$ denotes the conjugate transpose of a vector.
The following corollary presents a two-sided estimate for $\|E_{\rm TG}\|_{A}$.
Corollary 3.5.
Under the assumptions of Theorem 3.3, it holds that
(3.11) $\sqrt{1-\lambda_{n-r+s+1}(\widetilde{M}A)}\leq\|E_{\rm TG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(\widetilde{M}A)}.$
Proof.
Since $I-\Pi$ is an $L^{2}$-orthogonal projector of rank $n-s$, there exists a matrix with orthonormal columns, $V_{n-s}\in\mathbb{R}^{n\times(n-s)}$, such that
$I-\Pi=V_{n-s}V_{n-s}^{T}.$
According to the proof of Theorem 3.3, it follows that $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}V_{n-s}^{T}$ has $r-s$ positive eigenvalues and eigenvalue 00 with multiplicity $n-r+s$. Then
$\sigma_{\rm TG}$ $=\lambda_{n-r+s+1}\big{(}\widetilde{M}A(I-\Pi_{A})\big{)}$
$=\lambda_{n-r+s+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}(I-\Pi)\big{)}$
$=\lambda_{n-r+s+1}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}V_{n-s}^{T}\big{)}$
$=\lambda_{n-r+1}\big{(}V_{n-s}^{T}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}V_{n-s}\big{)}.$
Using Lemma 3.4, we obtain
$\lambda_{n-r+1}(\widetilde{M}A)\leq\sigma_{\rm TG}\leq\lambda_{n-r+s+1}(\widetilde{M}A),$
which, together with (3.7), yields the estimate (3.11). ∎
Remark 3.6.
From the relation
$I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}=\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)},$
we deduce that
$\lambda(\widetilde{M}A)=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset(-\infty,1].$
In addition,
$\lambda(\widetilde{M}A)$ $=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda\big{(}I-\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\subset[0,+\infty),$
where $\overline{M}$ is defined by (2.5). Hence, $\lambda(\widetilde{M}A)\subset[0,1]$.
4. An inexact variant of Algorithm 1 and its analysis
In this section, we analyze the convergence of an inexact variant of Algorithm 1, which is described by Algorithm 2. The output of its correction step, $\hat{\mathbf{e}}_{\rm c}$, is expected to be a good approximation to $\mathbf{e}_{\rm c}=A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$.
Algorithm 2  Inexact two-grid method.
1:Smoothing: $\mathbf{u}^{(1)}\leftarrow\mathbf{u}^{(0)}+M\big{(}\mathbf{f}-A\mathbf{u}^{(0)}\big{)}$
2:Restriction: $\mathbf{r}_{\rm c}\leftarrow P^{T}\big{(}\mathbf{f}-A\mathbf{u}^{(1)}\big{)}$
3:Coarse-grid correction: $\hat{\mathbf{e}}_{\rm c}\leftarrow\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket$triangleright $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket\approx A_{\rm c}^{\dagger}:\mathbb{R}^{n_{\rm c}}\rightarrow\mathbb{R}^{n_{\rm c}}$
4:Prolongation: $\mathbf{u}_{\rm ITG}\leftarrow\mathbf{u}^{(1)}+P\hat{\mathbf{e}}_{\rm c}$
This section is divided into two subsections. The first one focuses on the convergence analysis of Algorithm 2 with linear coarse solvers; the second one is devoted to the case of general coarse solvers.
4.1. The linear case
Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be an SPSD approximation to $A_{\rm c}$. In this subsection, we study the convergence of Algorithm 2 with $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket=B_{\rm c}^{\dagger}$.
We first prove an elementary lemma about the spectral relation between $A_{\rm c}$ and $B_{\rm c}$.
Lemma 4.1.
Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be SPSD. Then, there exist positive numbers $c_{1}$ and $c_{2}$ such that
(4.1) $c_{1}\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}\leq c_{2}\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}$
if and only if
(4.2) $\mathcal{R}(B_{\rm c})=\mathcal{R}(A_{\rm c}).$
Moreover, (4.1) is equivalent to the following inequality:
(4.3) $c_{2}^{-1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}B_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq c_{1}^{-1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}.$
Proof.
If (4.1) holds, then
$\mathcal{N}(B_{\rm c})=\mathcal{N}(A_{\rm c}),$
which is equivalent to (4.2). Conversely, if (4.2) holds, then
$\operatorname*{rank}(B_{\rm c})=\operatorname*{rank}(A_{\rm c})=s.$
Let
$A_{\rm c}=U_{1}\Sigma_{1}U_{1}^{T}\quad\text{and}\quad B_{\rm c}=U_{2}\Sigma_{2}U_{2}^{T},$
where, for each $i\in\{1,2\}$, $U_{i}\in\mathbb{R}^{n_{\rm c}\times s}$ has orthonormal columns and $\Sigma_{i}\succ 0$ is an $s\times s$ diagonal matrix. The relation (4.2) implies that
$\mathcal{R}(U_{2})=\mathcal{R}(U_{1}).$
Then, there exists an orthogonal matrix $\Omega_{s}\in\mathbb{R}^{s\times s}$ such that $U_{2}=U_{1}\Omega_{s}$. Hence,
$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c})}\frac{\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}}$ $=\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(U_{1}^{T})}\frac{\mathbf{v}_{\rm c}^{T}U_{1}\big{(}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\big{)}U_{1}^{T}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}U_{1}\Sigma_{1}U_{1}^{T}\mathbf{v}_{\rm c}}$
$=\min_{\mathbf{v}_{\rm s}\in\mathcal{R}(U_{1}^{T})\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\Sigma_{1}\mathbf{v}_{\rm s}}$
$=\min_{\mathbf{v}_{\rm s}\in\mathbb{R}^{s}\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\Sigma_{1}\mathbf{v}_{\rm s}}$
$=\lambda_{\min}\big{(}\Sigma_{1}^{-1}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\big{)}$
$=\lambda_{n_{\rm c}-s+1}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}.$
Also, it holds that
$\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c})}\frac{\mathbf{v}_{\rm c}^{T}B_{\rm c}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}\mathbf{v}_{\rm c}}=\lambda_{\max}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}.$
Therefore, the inequality (4.1) holds if $c_{1}\leq\lambda_{n_{\rm c}-s+1}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}$ and $c_{2}\geq\lambda_{\max}\big{(}A_{\rm c}^{\dagger}B_{\rm c}\big{)}$.
We next prove the equivalence between (4.1) and (4.3). In view of (4.1), we have
$U_{1}\big{(}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}-c_{1}\Sigma_{1}\big{)}U_{1}^{T}=B_{\rm c}-c_{1}A_{\rm c}\succeq 0,$
from which we deduce that $\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\succeq c_{1}\Sigma_{1}$ and thus $c_{1}^{-1}\Sigma_{1}^{-1}\succeq\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}$. Since
$c_{1}^{-1}A_{\rm c}^{\dagger}-B_{\rm c}^{\dagger}=U_{1}\big{(}c_{1}^{-1}\Sigma_{1}^{-1}-\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\big{)}U_{1}^{T},$
it follows that $c_{1}^{-1}A_{\rm c}^{\dagger}\succeq B_{\rm c}^{\dagger}$. Conversely, if $c_{1}^{-1}A_{\rm c}^{\dagger}\succeq B_{\rm c}^{\dagger}$, then $c_{1}^{-1}\Sigma_{1}^{-1}\succeq\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}$, which leads to $B_{\rm c}\succeq c_{1}A_{\rm c}$.
Similarly, one can show that $c_{2}A_{\rm c}\succeq B_{\rm c}$ if and only if $B_{\rm c}^{\dagger}\succeq c_{2}^{-1}A_{\rm c}^{\dagger}$. The proof is completed. ∎
Based on Lemma 4.1 and its proof, we can further derive the following spectral relation.
Lemma 4.2.
Let $B_{\rm c}\in\mathbb{R}^{n_{\rm c}\times n_{\rm c}}$ be an SPSD matrix satisfying (4.2), and let
(4.4) $\alpha_{1}=\lambda_{n_{\rm c}-s+1}\big{(}B_{\rm c}^{\dagger}A_{\rm c}\big{)}\quad\text{and}\quad\alpha_{2}=\lambda_{\max}\big{(}B_{\rm c}^{\dagger}A_{\rm c}\big{)}.$
Define
$\widetilde{B}_{\rm c}^{\dagger}:=2B_{\rm c}^{\dagger}-B_{\rm c}^{\dagger}A_{\rm c}B_{\rm c}^{\dagger}.$
If $\alpha_{2}<2$, then
(4.5) $\beta_{1}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\leq\beta_{2}\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}\quad\forall\,\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}},$
where
(4.6a) $\beta_{1}$ =$=\begin{cases}(2-\alpha_{1})\alpha_{1}&\text{if $\alpha_{2}\leq 1$},\\ \min\big{\{}(2-\alpha_{1})\alpha_{1},\,(2-\alpha_{2})\alpha_{2}\big{\}}&\text{if $\alpha_{1}\leq 1<\alpha_{2}<2$},\\ (2-\alpha_{2})\alpha_{2}&\text{if $1<\alpha_{1}\leq\alpha_{2}<2$},\end{cases}$
(4.6b) $\beta_{2}$ =$=\begin{cases}(2-\alpha_{2})\alpha_{2}&\text{if $\alpha_{2}\leq 1$},\\ 1&\text{if $\alpha_{1}\leq 1<\alpha_{2}<2$},\\ (2-\alpha_{1})\alpha_{1}&\text{if $1<\alpha_{1}\leq\alpha_{2}<2$}.\end{cases}$
Proof.
Recall that
$A_{\rm c}=U_{1}\Sigma_{1}U_{1}^{T}\quad\text{and}\quad B_{\rm c}=U_{1}\Omega_{s}\Sigma_{2}\Omega_{s}^{T}U_{1}^{T}.$
Then
$\widetilde{B}_{\rm c}^{\dagger}=U_{1}\Omega_{s}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\Omega_{s}^{T}U_{1}^{T}$
and hence
$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}$ $=\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(U_{1}^{T})}\frac{\mathbf{v}_{\rm c}^{T}U_{1}\Omega_{s}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\Omega_{s}^{T}U_{1}^{T}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}U_{1}\Omega_{s}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\Omega_{s}^{T}U_{1}^{T}\mathbf{v}_{\rm c}}$
$=\min_{\mathbf{v}_{\rm s}\in\mathcal{R}(\Omega_{s}^{T}U_{1}^{T})\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\mathbf{v}_{\rm s}}$
$=\min_{\mathbf{v}_{\rm s}\in\mathbb{R}^{s}\backslash\{0\}}\frac{\mathbf{v}_{\rm s}^{T}\big{(}2\Sigma_{2}^{-1}-\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\Sigma_{2}^{-1}\big{)}\mathbf{v}_{\rm s}}{\mathbf{v}_{\rm s}^{T}\big{(}\Omega_{s}^{T}\Sigma_{1}\Omega_{s}\big{)}^{-1}\mathbf{v}_{\rm s}}$
$=\lambda_{\min}\big{(}2\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}-\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}^{2}\big{)}.$
Similarly, it holds that
$\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\lambda_{\max}\big{(}2\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}-\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}^{2}\big{)}.$
Note that
$\lambda_{\min}\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}=\alpha_{1}\quad\text{and}\quad\lambda_{\max}\big{(}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\Sigma_{1}\big{)}=\alpha_{2},$
where $\alpha_{1}$ and $\alpha_{2}$ are given by (4.4). We then have
$\inf_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\beta_{1}\quad\text{and}\quad\sup_{\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}\backslash\mathcal{N}(A_{\rm c}^{\dagger})}\frac{\mathbf{v}_{\rm c}^{T}\widetilde{B}_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}{\mathbf{v}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{v}_{\rm c}}=\beta_{2},$
where $\beta_{1}$ and $\beta_{2}$ are given by (4.6a) and (4.6b), respectively. In addition,
$\widetilde{B}_{\rm c}^{\dagger}=U_{1}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}\big{(}2\Omega_{s}\Sigma_{2}\Omega_{s}^{T}-\Sigma_{1}\big{)}\Omega_{s}\Sigma_{2}^{-1}\Omega_{s}^{T}U_{1}^{T},$
which, together with the condition $\alpha_{2}<2$ (or, equivalently, $2\Omega_{s}\Sigma_{2}\Omega_{s}^{T}\succ\Sigma_{1}$), leads to
$\mathcal{N}\big{(}\widetilde{B}_{\rm c}^{\dagger}\big{)}=\mathcal{N}\big{(}U_{1}^{T}\big{)}=\mathcal{N}\big{(}A_{\rm c}^{\dagger}\big{)}.$
Thus, the inequality (4.5) holds for all $\mathbf{v}_{\rm c}\in\mathbb{R}^{n_{\rm c}}$. ∎
Remark 4.3.
In the extreme case $B_{\rm c}=A_{\rm c}$, one has $\alpha_{1}=\alpha_{2}=1$ and $\beta_{1}=\beta_{2}=1$. We remark that $\alpha_{2}<2$ is a mild condition, which can be easily satisfied (e.g., one may simply scale $B_{\rm c}$ if necessary).
It is easy to see that the iteration matrix of Algorithm 2 with $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket=B_{\rm c}^{\dagger}$ is
$E_{\rm ITG}=\big{(}I-PB_{\rm c}^{\dagger}P^{T}A\big{)}(I-MA),$
which satisfies
$\mathbf{u}-\mathbf{u}_{\rm ITG}=E_{\rm ITG}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}.$
Due to
$A^{\frac{1}{2}}E_{\rm ITG}=\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}A^{\frac{1}{2}},$
it follows from Lemma 2.1 that
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\|E_{\rm ITG}\|_{A}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}$
with
$\|E_{\rm ITG}\|_{A}=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\big{\|}\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\big{\|}_{2}}{\|\mathbf{v}\|_{2}}.$
Then
$\|E_{\rm ITG}\|_{A}^{2}$ $=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}PB_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}^{2}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
$=\max_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}$
$=1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}},$
where
(4.7) $F_{\rm ITG}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
Since $A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\succeq 0$ and $\widetilde{B}_{\rm c}^{\dagger}\succeq 0$, it follows that $F_{\rm ITG}\succeq 0$. For any $\mathbf{v}\in\mathcal{N}(F_{\rm ITG})$, we have
$0=\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}=\mathbf{v}^{T}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\mathbf{v}+\mathbf{v}^{T}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}A^{\frac{1}{2}}P\widetilde{B}_{\rm c}^{\dagger}P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v},$
from which we deduce that
$\mathbf{v}\in\mathcal{N}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}\quad\text{and}\quad P^{T}A^{\frac{1}{2}}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\mathbf{v}\in\mathcal{N}\big{(}\widetilde{B}_{\rm c}^{\dagger}\big{)}=\mathcal{N}\big{(}A_{\rm c}^{\dagger}\big{)}.$
In view of (2.9), we have $\mathbf{v}\in\mathcal{N}(F_{\rm TG})$, which yields
$\mathcal{N}(F_{\rm ITG})\subseteq\mathcal{N}(F_{\rm TG}).$
Similarly, one can prove that $\mathcal{N}(F_{\rm TG})\subseteq\mathcal{N}(F_{\rm ITG})$. Thus,
(4.8) $\mathcal{N}(F_{\rm ITG})=\mathcal{N}(F_{\rm TG}),$
which implies that $\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}>0$ if and only if $\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0$.
Note that
$\|E_{\rm ITG}\|_{A}<1$ $\Longleftrightarrow\mathbf{v}^{T}F_{\rm ITG}\mathbf{v}>0\quad\forall\,\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$
$\Longleftrightarrow\mathbf{v}^{T}F_{\rm TG}\mathbf{v}>0\quad\forall\,\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}$
$\Longleftrightarrow\|E_{\rm TG}\|_{A}<1.$
Hence, (2.10) is a sufficient and necessary condition for $\|E_{\rm ITG}\|_{A}<1$. In addition, from (3.4) and (4.8), we have
$\mathcal{N}(F_{\rm ITG})=\mathcal{N}(A).$
Similarly to the proof of Lemma 3.1, one can show that
(4.9) $\|E_{\rm ITG}\|_{A}=\sqrt{1-\lambda_{n-r+1}(F_{\rm ITG})}.$
Let
$F_{i}=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{i}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}$
for $i=1,2$. Using (4.5), (4.7), and (4.9), we obtain
(4.10) $\sqrt{1-\lambda_{n-r+1}(F_{2})}\leq\|E_{\rm ITG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(F_{1})}.$
To estimate the eigenvalues $\lambda_{n-r+1}(F_{1})$ and $\lambda_{n-r+1}(F_{2})$, we need the following eigenvalue inequalities; see, e.g., [12, Theorem 4.3.1].
Lemma 4.4.
Let $H_{1}\in\mathbb{C}^{n\times n}$ and $H_{2}\in\mathbb{C}^{n\times n}$ be Hermitian. For each $k=1,\ldots,n$, it holds that
(4.11a) $\lambda_{k}(H_{1}+H_{2})$ $\geq\lambda_{k-j+1}(H_{1})+\lambda_{j}(H_{2})\quad\forall\,j=1,\ldots,k,$
(4.11b) $\lambda_{k}(H_{1}+H_{2})$ $\leq\lambda_{k+\ell}(H_{1})+\lambda_{n-\ell}(H_{2})\quad\forall\,\ell=0,\ldots,n-k.$
In addition to (4.11a) and (4.11b), we need the following technical lemma.
Lemma 4.5.
Let $\Pi$, $\widetilde{M}$, and $\Pi_{A}$ be defined by (2.7), (3.5), and (3.6), respectively. Then
(4.12a) $\lambda_{\min}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=0,$
(4.12b) $\lambda_{\max}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=1-\delta_{\rm TG},$
where
(4.13) $\delta_{\rm TG}=\begin{cases}\lambda_{n-s+1}(\widetilde{M}A\Pi_{A})&\text{if $\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}=\{0\}$},\\[2.0pt] 0&\text{otherwise}.\end{cases}$
Proof.
The identity (4.12a) is obvious, because $\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\succeq 0$ is singular. Next, we focus on the proof of (4.12b).
Since $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\succeq 0$ and $\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset[0,1]$, it follows from (3.10) that $X_{1}\succeq 0$ and $\lambda(X_{1})\subset[0,1]$. By (3.9) and (3.10), we have
$\big{(}I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\Pi=Q_{2}\begin{pmatrix}I_{s}-X_{1}&0\\ -X_{2}^{T}&0\end{pmatrix}Q_{2}^{T}.$
Then
(4.14) $\lambda_{\max}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=\lambda_{\max}\big{(}\big{(}I-A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\Pi\big{)}$
$=1-\lambda_{\min}(X_{1}).$
Direct computation yields
$\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi=Q_{2}\begin{pmatrix}X_{1}&0\\ 0&0\end{pmatrix}Q_{2}^{T},$
where we have used the relations (3.9) and (3.10). It follows that
(4.15) $\operatorname*{rank}(X_{1})=\operatorname*{rank}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}.$
We then prove that $\operatorname*{rank}(X_{1})=s$ (i.e., $X_{1}\succ 0$) if and only if
(4.16) $\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\cap\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}=\{0\}.$
  • Recall that
    $\mathcal{R}(\Pi)=\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}\quad\text{and}\quad\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
    If (4.16) holds, then
    $\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}\subseteq\mathcal{N}(\Pi),$
    which, combined with the fact $\mathcal{N}(\Pi)\subseteq\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}$, leads to
    $\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}=\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
    Hence,
    $\operatorname*{rank}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}=\operatorname*{rank}\big{(}P^{T}A^{\frac{1}{2}}\big{)}=\operatorname*{rank}(A_{\rm c})=s,$
    which, together with (4.15), leads to $\operatorname*{rank}(X_{1})=s$.
  • Conversely, if $\operatorname*{rank}(X_{1})=s$, then (4.16) holds. Otherwise, there must exist a vector $\mathbf{v}\in\mathbb{R}^{n}\backslash\mathcal{N}(\Pi)$ such that
    $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\mathbf{v}=0,$
    i.e., $\mathbf{v}\in\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}\backslash\mathcal{N}(\Pi)$. This contradicts with the fact
    $\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}=\mathcal{N}(\Pi),$
    which follows from the relations $\mathcal{N}(\Pi)\subseteq\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}$ and
    $\operatorname*{rank}(\Pi)=s=\operatorname*{rank}(X_{1})=\operatorname*{rank}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}.$
The above analysis shows that
[…]
Note that
(4.18) $\lambda\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}=\lambda\big{(}\widetilde{M}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}\big{)}=\lambda(\widetilde{M}A\Pi_{A}).$
Combining (4.14), (4.17), and (4.18), we can arrive at the identity (4.12b). ∎
Remark 4.6.
The relation (4.16) will hold if $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$, since
$\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\widetilde{M}^{\frac{1}{2}}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A)\quad\text{and}\quad\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}\subseteq\mathcal{R}(A).$
Therefore, a sufficient condition for (4.16) to hold is $\widetilde{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, as discussed in Remark 2.6.
We are now ready to present a two-sided estimate for $\|E_{\rm ITG}\|_{A}$.
Theorem 4.7.
Under the assumptions of Algorithm 2 and Lemma 4.2, the inexact two-grid convergence factor $\|E_{\rm ITG}\|_{A}$ satisfies
(4.19) $\mathscr{L}(\beta_{2})\leq\|E_{\rm ITG}\|_{A}\leq\mathscr{U}(\beta_{1})$
with
$\mathscr{L}(\beta_{2})=\sqrt{1-\min\big{\{}\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}),\,\sigma_{\rm TG}\big{\}}}$
and
(4.20) $\mathscr{U}(\beta_{1})=\sqrt{1-\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,\beta_{1}\sigma_{\rm TG},\,\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG})\big{\}}},$
where $\widetilde{M}$, $\sigma_{\rm TG}$, $\beta_{1}$, $\beta_{2}$, and $\delta_{\rm TG}$ are given by (3.5), (3.8), (4.6a), (4.6b), and (4.13), respectively.
Proof.
In view of (4.10), one can get a two-sided estimate for $\|E_{\rm ITG}\|_{A}$ by bounding the eigenvalues $\lambda_{n-r+1}(F_{1})$ and $\lambda_{n-r+1}(F_{2})$.
A lower bound for $\lambda_{n-r+1}(F_{1})$: By (4.11a), we have
$\lambda_{n-r+1}(F_{1})$ $\geq\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{1}\lambda_{1}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n-r+1}(\widetilde{M}A),$
where we have used the facts (4.12a) and $\lambda\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}$. In addition, we have
$\lambda_{n-r+1}(F_{1})$ $=\lambda_{n-r+1}\big{(}(1-\beta_{1})A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{1}F_{\rm TG}\big{)}$
$\geq(1-\beta_{1})\lambda_{1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{1}\lambda_{n-r+1}(F_{\rm TG})$
$=\beta_{1}\sigma_{\rm TG}$
and
$\lambda_{n-r+1}(F_{1})$ $=\lambda_{n-r+1}\big{(}F_{\rm TG}-(1-\beta_{1})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\geq\lambda_{n-r+1}(F_{\rm TG})-(1-\beta_{1})\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG}),$
where, in the last equality, we have used the fact (4.12b). Hence,
(4.21) $\lambda_{n-r+1}(F_{1})\geq\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,\beta_{1}\sigma_{\rm TG},\,\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG})\big{\}}.$
An upper bound for $\lambda_{n-r+1}(F_{2})$: Using (4.11b) and (4.12b), we obtain
$\lambda_{n-r+1}(F_{2})$ $=\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{2}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\leq\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{2}\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}).$
Moreover,
$\lambda_{n-r+1}(F_{2})$ $=\lambda_{n-r+1}\big{(}F_{\rm TG}-(1-\beta_{2})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\leq\lambda_{n-r+1}(F_{\rm TG})-(1-\beta_{2})\lambda_{1}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\sigma_{\rm TG},$
where, in the last equality, we have used the fact (4.12a). Thus,
(4.22) $\lambda_{n-r+1}(F_{2})\leq\min\big{\{}\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}),\,\sigma_{\rm TG}\big{\}}.$
The estimate (4.19) then follows by combining (4.10), (4.21), and (4.22). ∎
Remark 4.8.
By (4.11a) and (4.12b), we have
$1-\delta_{\rm TG}$ $=\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n}\big{(}F_{\rm TG}-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
$\geq\lambda_{n-r+1}(F_{\rm TG})-\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
$=\sigma_{\rm TG}-\lambda_{n-r+1}(\widetilde{M}A),$
which gives
$\sigma_{\rm TG}\leq 1-\delta_{\rm TG}+\lambda_{n-r+1}(\widetilde{M}A).$
According to (4.6a) and (4.6b), we deduce that $0<\beta_{1}\leq\beta_{2}\leq 1$.
  • (i)
    If $\beta_{2}\rightarrow 0^{+}$, then $F_{i}\rightarrow A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}$ for $i=1,2$, which, together with (4.10), yields
    $\|E_{\rm ITG}\|_{A}\rightarrow\sqrt{1-\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}}=\sqrt{1-\lambda_{n-r+1}(\widetilde{M}A)}.$
    This asymptotic behavior can also be observed from (4.19) as $\beta_{2}\rightarrow 0^{+}$.
  • (ii)
    In the special case $B_{\rm c}=A_{\rm c}$ (i.e., $\beta_{1}=\beta_{2}=1$), we have
    $\mathscr{L}(1)=\mathscr{U}(1)=\sqrt{1-\sigma_{\rm TG}},$
    which leads to the convergence identity (3.7).
Remark 4.9.
Since
$\lambda(\widetilde{M}A)=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset[0,1]\quad\text{and}\quad\operatorname*{rank}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\leq\operatorname*{rank}\big{(}A^{\frac{1}{2}}\big{)}=r,$
it follows that
$\lambda_{n-r+1}(\widetilde{M}A)>0\Longleftrightarrow\operatorname*{rank}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}=r\Longleftrightarrow\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A),$
which will be satisfied if $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$.
4.2. The general case
In this subsection, we present a convergence estimate for Algorithm 2 with general coarse solver $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket$, as described in the following theorem.
Theorem 4.10.
Under the assumptions of Algorithm 2, if
(4.23) $\big{\|}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}\leq\varepsilon\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}$
for some $\varepsilon\in[0,1)$, then
(4.24) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\mathscr{U}(1-\varepsilon^{2})\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A},$
where $\mathscr{U}(\cdot)$ is given by (4.20).
Proof.
The condition (4.23) reads
$2\big{(}A_{\rm c}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}\big{)}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}.$
Note that $A_{\rm c}A_{\rm c}^{\dagger}$ is a projector onto $\mathcal{R}(A_{\rm c})$ and
$\mathbf{r}_{\rm c}=P^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}\in\mathcal{R}\big{(}P^{T}A^{\frac{1}{2}}\big{)}=\mathcal{R}(A_{\rm c}).$
We then have
(4.25) $2\mathbf{r}_{\rm c}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}.$
From Algorithm 2, we have
$\mathbf{u}-\mathbf{u}_{\rm ITG}=\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket.$
Then
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}$ $=\big{(}\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}$
$=\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-2\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}AP\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}$
$=\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-2\mathbf{r}_{\rm c}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}$
$\leq\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}-(1-\varepsilon^{2})\mathbf{r}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}-(1-\varepsilon^{2})\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A^{\frac{1}{2}}\big{(}I-(1-\varepsilon^{2})\Pi\big{)}A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)},$
where we have used the inequality (4.25). Due to
$\mathbf{u}-\mathbf{u}^{(1)}=(I-MA)\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
and
$\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}=I-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}},$
it follows that
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}^{T}A^{\frac{1}{2}}(I-F)A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
with
$F=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+(1-\varepsilon^{2})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
Then
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\bigg{(}1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}\bigg{)}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}^{2}.$
Similarly to the proof of Lemma 3.1, one can prove that
$\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}=\lambda_{n-r+1}(F).$
Thus,
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(F)}\,\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}.$
The estimate (4.24) can be obtained by bounding $\lambda_{n-r+1}(F)$ from below. Indeed, it holds that
$\lambda_{n-r+1}(F)\geq\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,(1-\varepsilon^{2})\sigma_{\rm TG},\,\sigma_{\rm TG}-\varepsilon^{2}(1-\delta_{\rm TG})\big{\}};$
see the proof of (4.21). This completes the proof. ∎
Remark 4.11.
From Algorithms 1 and 2, we have
$\|\mathbf{u}_{\rm TG}-\mathbf{u}_{\rm ITG}\|_{A}=\big{\|}P\big{(}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}\big{\|}_{A}=\big{\|}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}},$
i.e., the perturbation bound of two-grid solution is equal to that of coarse solution. Therefore, the accuracy of two-grid solver can be controlled by that of coarse solver.
5. Conclusions
Two-grid convergence theory plays an important role in the design and analysis of multigrid methods. In this paper, we have established a new convergence theory for two-grid methods applied to SPSD linear systems. In the case where the Moore–Penrose inverse of coarse-grid matrix is used as a coarse solver, we derive a succinct identity for the convergence factor of two-grid methods. More generally, we present some convergence estimates for two-grid methods with approximate coarse solvers (both linear and general types are considered). A distinctive feature in comparison with the existing ones is that our theory does not require any additional assumptions on the system matrix, especially on its null space.
Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (Grant No. 12401479), the Natural Science Foundation of Jiangsu Province (Grant No. BK20241257), and the Scientific Research Foundation of the Southeast University (Grant No. 4007012324).